OpenStudy (anonymous):

What is the equation of the line in point-slope form that passes through the given point and has the given slope? point; (8,-8) and slope;3

OpenStudy (anonymous):

y - y1 = m(x - x1) m, your slope is 3 (8,-8) x1 = 8 and y1 = -8 now what can you do with that ?

OpenStudy (anonymous):

just sub in your info into the formula

OpenStudy (anonymous):

okay and then what?

OpenStudy (anonymous):

that is your answer.

OpenStudy (anonymous):

I got y-8=3(x-8) is that right?

OpenStudy (jdoe0001):

\(\bf (x_1,y_1) \implies (8, -8)\\\large \\\quad \\ \begin{array}{llll} (&\bf 8, &\bf -8)\\ y-&\bf y_1=m(x-&\bf x_1) \end{array}\)

OpenStudy (jdoe0001):

ahemm... even ... wrong there.. shoot

OpenStudy (anonymous):

pay attention to the signs.. y - (-8) = 3(x - 8) which is : y + 8 = 3(x - 8)

OpenStudy (jdoe0001):

\(\bf (x_1,y_1) \implies (8, -8)\\\large \\\quad \\ \begin{array}{llll} (&\bf -8, &\bf 8)\\ y-&\bf y_1=m(x-&\bf x_1) \end{array}\)